# Accelerating ViT Inference on FPGA through Static and Dynamic Pruning

### 21 Mar 2024 v1

@@ IV-B Dynamic Token Pruning
Dynamic Token Pruning prunes along the token dimension $N$. The redundancy along the token dimension comes from the fact that several patches within an input image are inattentive, contributing insignificantly to the final learned model [35, 34, 28, 36]. Since ViTs can inherently handle inputs with arbitrary number of tokens (patches), we exploit this independence of the input token dimension from the model parameter dimension(s) by dropping inattentive tokens. Specifically, to classify tokens into attentive and inattentive tokens, we use a non-parametric approach [28]. The attention $\mathbf{A}$ computed within the MSA (Equation 3) is utilized to perform attentive token identification. In MSA, the attention score $\mathbf{A}_{h}$ is generated by each head. We aggregate the above score vector across all the heads using $\mathbf{\mathcal{S}}=\frac{1}{H}\sum_{h}\mathbf{A}_{h}\text{ }\mbox{where}\text{ }\mathbf{\mathcal{S}}\in\mathbb{R}^{N}$ where $\mathbf{\mathcal{S}}$ represents the importance score of every single token. Based on a keep-rate $r_{t}$, a total of $\left\lceil(N-1)r_{t}\right\rceil$ tokens with the top scores in $\mathbf{\mathcal{S}}$ are retained. The remaining inattentive tokens are fused into a single token by performing a weighted aggregation of these tokens with respect to their respective scores in $\mathbf{\mathcal{S}}$. The above token dropping is performed via a token dropping module (TDM) inserted between the MSA and the MLP modules (Figure 4), with the tokens dropped dynamically during both training and inference.
@@ IV-C Simultaneous Pruning
In order to recove accuracy for the pruned model, we utilize the knowledge distillation technique commonly used to transfer knowledge from an already trained larger teacher model to a smaller student model [42]. The class logits associated with the teacher and the student networks are used to compute a distillation loss at a distillation temperature $T$:
@@ IV-D Computational Complexity: Pruned Model
We analyze the computational complexity for the proposed pruned model. The complexity of an encoder is described in table II. $\alpha$ is the average ratio of retained weight blocks to the total weight blocks (retained and pruned) within a column of blocks in parameter matrices $\mathbf{W}_{p}$ (computed after the removal of heads pruned in their entirety), $\alpha^{\prime}$ is defined similarly, but for matrix $\mathbf{W}_{\text{proj}}$. $H_{\text{kept}}$ are the number of heads retained within MSA, $N_{\text{kept}}$ are the total retained tokens after token dropping ($\approx Nr_{t}$). $\alpha^{\text{mlp}}$ is the ratio of retained neurons (same for both $\mathbf{W}_{\text{int}}$ and $\mathbf{W}_{\text{out}}$). Note that $\alpha^{\text{mlp}}=r_{b}$.
@@ V Hardware Design
In this Section, we introduce our hardware design to accelerate the pruned model on FPGA platform. To be specific: in Section V-A, we introduce the data format and layout that stores the sparse (and dense) weight matrices and input data; in Section V-B, we introduce the main components in the proposed hardware architecture. In Section V-C, we introduce the workflow for executing the pruned ViT encoder using the proposed hardware design.
@@ V-A Data Format and Layout
Due to structured block pruning, all weight and feature matrices are partitioned into data blocks of same size $b\times b$, where $b$ is the block size in block pruning. All the data in the same blocks are stored contiguously. The dense matrix are stored in block-wise row-major order such that all the data blocks of the same row are stored contiguous in memory space. The weight matrices are stored in column-major order such that all the unpruned data blocks in the same column are stored contiguous in memory space as shown in Figure 5. Note that for the sparse weight matrices, only unpruned data blocks are stored. For each column in a block-wise sparse weight matrix, we include a header at its beginning that encodes row indices of the present blocks and the length of the column block. For simplicity, in the rest of the paper, a row of matrix denotes a row of data blocks and a column of matrix denotes a column of data blocks.
@@ V-B Hardware Overview
As shown in Figure 6, the architecture design comprises of: (1) Multi-level Parallelism Compute Array (MPCA), (2) Element-wise Module (EM), (3) Token Dropping Hardware Module (TDHM). Besides, there are on-chip buffers, including Global Feature Buffer (GFB) that stores the feature matrices, (2) Column Buffers (CB) that store the weight matrix, (3) Result Buffers (RB) that store the results of the current layer.
In MPCA, the computation units are organized into multiple levels. A MPCA has $p_{\text{h}}$ parallel Computing Head Modules (CHMs). Each CHM has a 2-D array of Processing Elements (PEs) of size $p_{\text{t}}\times p_{\text{c}}$. Each Processing Element has an array of computation units of size $p_{\text{pe}}\times p_{\text{pe}}$. Essentially, $p_{\text{h}}$, $p_{\text{t}}$, $p_{\text{c}}$, $p_{\text{pe}}^{2}$ are the computation parallelism in the head dimension, input token dimension, weight column dimension, and the data parallelism within the data blocks, respectively. The Element-wise Module (EM) performs element-wise GELU and exponentiation. The Token Dropping Hardware Module (TDHM) performs dynamical token dropping (Section IV-B).
The computation units in MPCA are organized to multi-level, because (1) it enables massive data parallelism in MSA and MLP, (2) it enables data reuse/sharing within CHM. The PEs in the same column of CHM can share the same weight block while the PEs in the same rows of CHM can share the same input token block. This data sharing strategy simplifies the compute complicated by the irregular data access pattern of block-wise weight pruning. (3) Through selecting proper $p_{\text{c}}$ with load balancing strategy, we can alleviate the load imbalance caused due to block-wise weight pruning.
@@ V-C Workflow
The proposed accelerator executes the input model layer-by-layer with each layer executed using the same set of computational resources (e.g., MPCA). Eeach layer as follows:
@@ V-C1 MSA Execution
MSA is divided into four stages (shown in Figure 7): stage (i) computes $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ through $[\mathbf{Q},\mathbf{K},\mathbf{V}]=\mathbf{Z}[\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}]$. $\mathbf{Z}$ is the input token matrix and [][\;][ ] is matrix concatenation. Let $\mathbf{Q}_{h},\mathbf{K}_{h}$ and $\mathbf{V}_{h}$ denote the query, key and value matrices for a head $h$, where $(0\leq h<H)$. The algorithm for executing each matrix multiplication (e.g., dense token matrix $\mathbf{Z}$ multiply by sparse weight matrices $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$) in MSA using MPCA is shown in Algorithm 2 and an example is shown in Figure 8. Each CHM computes its corresponding head $[\mathbf{Q}_{h},\mathbf{K}_{h},\mathbf{V}_{h}]$. Since we perform block partitioning for each matrix, the computation is executed in a block-wise fashion. Within each CHM, the PEs within a column share the same column of weight, which is stored in the column buffer (CB). PEs of the same row share the same row of $\mathbf{Z}$. We use $\text{PE}(i,j,k)$ to denote a PE in the $k^{\text{th}}$ CHM at location $(i,j)$ within the CHM. A column of PEs: $\text{PE}(:,j,k)$ share the same column of weight with the corresponding header information (See Figure 5). Thus, each PE in $j^{\text{th}}$ column utilizes the shared header indices (Figure 5) to fetch the corresponding data block from the token matrix (in the local input buffer) to perform block-wise matrix multiplication. The partial results are accumulated in local result buffers.
The stage (ii) computes the attention scores $\mathbf{A}_{h}=\mbox{softmax}(\frac{\mathbf{Q}_{h}\mathbf{K_{h}}^{T}}{\sqrt{D^{\prime}}}),\text{ }(0\leq i<H)$. $\mathbf{Q}_{h}\mathbf{K}_{h}^{T}$ is dense block-wise matrix multiplication executed via the MPCA module (See Algorithm 2) with $\mathbf{Q}$ being buffered in the GFB and $\mathbf{K}^{T}$ being buffered in the CB. The output data blocks of $\mathbf{Q}_{h}\mathbf{K}_{h}^{T}$ are sent to EM module for element-wise scaling (by $1/\sqrt{D^{\prime}}$) and exponentiation to obtain $\text{exp}(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{D^{\prime}}})$. Then, we utilize MPCA to compute the scaling factors for $\mbox{softmax}(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{D^{\prime}}})$. The rows of matrix $\text{exp}(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{D^{\prime}}})$ and their corresponding computed scaling factors, are streamed from MPCA to EM to perform the scaling to obtain the attention scores, $\mathbf{A}^{H}$.
@@ V-C2 MLP Execution
1:Input matrix $\mathbf{X}\in\mathbb{R}^{M_{1}\times M_{2}}$; weight matrix $\mathbf{W}=[\mathbf{W}_{0},\mathbf{W}_{1},...,\mathbf{W}_{H-1}]\in\mathbb{R}^{M_{2}\times D}$, where each $\mathbf{W}_{h}\in\mathbb{R}^{M_{2}\times D^{\prime}}$ ($0\leq h<H$); $D=HD^{\prime}$ where $H$ denotes the number of heads and $D^{\prime}$ is the dimension per head; block size $b$;
@@
7:for $i=0\;\text{ to }\;\left\lceil\frac{H}{p_{\text{h}}}\right\rceil-1$ do
8:    for each $\text{CHM}_{j}$ with $j=0\;\text{ to }\;p_{\text{h}}-1$ Parallel do
9:         // $\text{CHM}_{j}$ computes $\mathbf{Y}_{j+ip_{\text{h}}}$
10:         for $k=0\;\text{ to }\;\left\lceil\left\lceil\frac{D^{\prime}}{b}\right\rceil/{p_{\text{c}}}\right\rceil-1$ do
11:             // Load weights into CB
12:             for $l=0\;\text{ to }\;\left\lceil\left\lceil\frac{M}{b}\right\rceil/p_{\text{t}}\right\rceil-1$ do
13:                 // Load data (partition of $\mathbf{X}$) into GFB
14:                 for each $\text{PE}_{j}(m,n)$ in $\text{CHM}_{j}$ Parallel do
15:                     // $\text{PE}_{j}(m,n)$ computes output block                             $\mathbf{Y}_{j+ip_{\text{h}}}[m+lp_{\text{t}},n+kp_{\text{c}}]$
16:                     if MPCA mode is SBMM then
17:                         Fetch $\mathbf{X}[m+lp_{\text{t}},\text{idx}]$ from GFB for all idx                                in the header of $\mathbf{W}_{j+ip_{\text{h}}}[:,n+kp_{\text{c}}]$
18:                         Compute $\mathbf{Y}_{j+ip_{\text{h}}}[m+lp_{\text{t}},n+kp_{\text{c}}]$ using                                the fetched input blocks from GFB
19:                     else
20:                         // MPCA mode is DBMM
21:                         Fetch $\mathbf{X}[m+lp_{\text{t}},:]$ from GFB
22:                         Compute $\mathbf{Y}_{j+ip_{\text{h}}}[m+lp_{\text{t}},n+kp_{\text{c}}]$ using                                all the input blocks from GFB
@@ V-E1 Resource Consumption Model
We perform theoretical analysis for the performance achieved by the codesign and its hardware resource utilization. We denote the total computational resources utilized by the MPCA, EM and TDHM as $R_{\text{MPCA}},R_{\text{EM}}$ and $R_{\text{TDHM}}$, respectively. $R_{\text{MPCA}}$ is proportional to the total number of computation units: $p_{\text{t}}p_{\text{h}}p_{\text{c}}p_{\text{pe}^{2}}$. Compared to $R_{\text{MPCA}}$, the resources used by $R_{\text{TDHM}}$ and $R_{\text{EM}}$ are negligible, and thus ignored for analysis. The total $R_{\text{Total}}$ (DSPs and LUTs) are $R_{\text{Total}}=(c_{1}p_{\text{t}}p_{\text{h}}p_{\text{c}}p_{\text{pe}}^{2},c_{2}p_{\text{t}}p_{\text{h}}p_{\text{c}}p_{\text{pe}}^{2})$, where $c_{1}$ and $c_{2}$ denote the amount of DSPs and LUTs utilized by a single computation unit. The size of the (global) feature buffer, column buffer and the (global) result buffer, associated with the MPCA, are $b^{2}p_{t}\gamma$, $b^{2}p_{c}\gamma$ and $b^{2}p_{\text{t}}p_{\text{h}}p_{\text{c}}$, respectively. Here, $b$ is the block size and $\gamma$ is a constant that equals the (maximum) total number of row blocks required to compute a single output block. We match the buffer sizes across compute units to improve the dataflow performance of the accelerator. This gives a total buffer size of $4\times\max({b^{2}p_{\text{t}}p_{\text{h}}p_{\text{c}},b^{2}p_{t}\gamma})$ for the EM module and $2\times\max({b^{2}p_{\text{t}}p_{\text{h}}p_{\text{c}},b^{2}p_{t}\gamma})$ for the TDHM module. The EM module requires a buffer to store the input, scaling factor, addtion factor and the output. similarly, TDHM requires an input and an output buffer. The total size of required buffers $B_{\text{Total}}$ is given as, $B_{\text{Total}}=b^{2}p_{t}\gamma+b^{2}p_{c}\gamma+b^{2}p_{\text{t}}p_{\text{h}}p_{\text{c}}+6\times\max({b^{2}p_{\text{t}}p_{\text{h}}p_{\text{c}},b^{2}p_{t}\gamma})$. $R_{\text{Total}}$ and $B_{\text{Total}}$ are the estimation of resource utilizations. The main design $p_{\text{t}}$, $p_{\text{h}}$ and $p_{\text{c}}$, with $\gamma$ are empirically set according the resource of target FPGA platform (See Section VI for details).

### 12 Apr 2024 v2

@@ IV-B Dynamic Token Pruning
Dynamic Token Pruning prunes along the token dimension $N$. The redundancy along the token dimension comes from the fact that several patches within an input image are inattentive, contributing insignificantly to the final learned model [35, 34, 28, 36]. Since ViTs can inherently handle inputs with an arbitrary number of tokens (patches), we exploit this independence of the input token dimension from the model parameter dimension(s) by dropping inattentive tokens. Specifically, to classify tokens into attentive and inattentive tokens, we use a non-parametric approach [28]. The attention $\mathbf{A}$ computed within the MSA (Equation 3) is utilized to perform attentive token identification. In MSA, the attention score $\mathbf{A}_{h}$ is generated by each head. We aggregate the above score vector across all the heads using $\mathbf{\mathcal{S}}=\frac{1}{H}\sum_{h}\mathbf{A}_{h}\text{ }\mbox{where}\text{ }\mathbf{\mathcal{S}}\in\mathbb{R}^{N}$ and represents the importance score of every single token. Based on a keep-rate $r_{t}$, a total of $\left\lceil(N-1)r_{t}\right\rceil$ tokens with the top scores in $\mathbf{\mathcal{S}}$ are retained. The remaining inattentive tokens are fused into a single token by performing a weighted aggregation of these tokens with respect to their respective scores in $\mathbf{\mathcal{S}}$. The above token dropping is performed via a token dropping module (TDM) inserted between the MSA and the MLP modules (Figure 4), with the tokens dropped dynamically during both training and inference.
@@ IV-C Simultaneous Pruning
To recover accuracy for the pruned model, we utilize the knowledge distillation technique commonly used to transfer knowledge from an already trained larger teacher model to a smaller student model [43]. The class logits associated with the teacher and the student networks are used to compute a distillation loss at a distillation temperature $T$:
@@ IV-D Computational Complexity: Pruned Model
We analyze the computational complexity for the proposed pruned model. The complexity of an encoder is described in table II. $\alpha$ is the average ratio of retained weight blocks to the total weight blocks (retained and pruned) within a column of blocks in parameter matrices $\mathbf{W}_{p}$ (computed after the removal of heads pruned in their entirety). $\alpha^{\prime}$ is defined similarly, but for matrix $\mathbf{W}_{\text{proj}}$. $H_{\text{kept}}$ are the number of heads retained within MSA. $N_{\text{kept}}$ are the total retained tokens after token dropping ($\approx Nr_{t}$). $\alpha^{\text{mlp}}$ is the ratio of retained neurons (same for both $\mathbf{W}_{\text{int}}$ and $\mathbf{W}_{\text{out}}$). Note that $\alpha^{\text{mlp}}=r_{b}$.
@@ V Hardware Design
In this Section, we introduce our hardware design to accelerate the pruned model on the FPGA platform. To be specific: in Section V-A, we introduce the data format and layout that store the sparse (and dense) weight matrices and input data; in Section V-B, we introduce the main components in the proposed hardware architecture. In Section V-C, we introduce the workflow for executing the pruned ViT encoder using the proposed hardware design.
@@ V-A Data Format and Layout
Due to structured block pruning, all weight and feature matrices are partitioned into data blocks of the same size $b\times b$. All the data in the same blocks are stored contiguously. The dense matrix is stored in block-wise row-major order such that all the data blocks of the same row are stored contiguously in memory space. The weight matrices are stored in column-major order such that all the unpruned data blocks in the same column are stored contiguous in memory space as shown in Figure 5. Note that for the sparse weight matrices, only unpruned data blocks are stored. For each column in a block-wise sparse weight matrix, we include a header at its beginning that encodes row indices of the present blocks and the length of the column block. For simplicity, in the rest of the paper, a row of matrix denotes a row of data blocks, and a column of matrix denotes a column of data blocks.
@@ V-B Hardware Overview
As shown in Figure 6, the architecture design comprises of: (1) Multi-level Parallelism Compute Array (MPCA), (2) Element-wise Module (EM), (3) Token Dropping Hardware Module (TDHM). Besides, there are on-chip buffers, including a Global Feature Buffer (GFB) that stores the feature matrices, (2) Column Buffers (CB) that store the weight matrix, (3) Result Buffers (RB) that store the results of the current layer.
In MPCA, the computation units are organized into multiple levels. An MPCA has $p_{\text{h}}$ parallel Computing Head Modules (CHMs). Each CHM has a 2-D array of Processing Elements (PEs) of size $p_{\text{t}}\times p_{\text{c}}$. Each Processing Element has an array of computation units of size $p_{\text{pe}}\times p_{\text{pe}}$. Essentially, $p_{\text{h}}$, $p_{\text{t}}$, $p_{\text{c}}$, $p_{\text{pe}}^{2}$ are the computation parallelism in the head dimension, input token dimension, weight column dimension, and the data parallelism within the data blocks, respectively. The Element-wise Module (EM) performs element-wise GELU and exponentiation. The Token Dropping Hardware Module (TDHM) performs dynamical token dropping (Section IV-B).
The computation units in MPCA are organized to multi-level because (1) it enables massive data parallelism in MSA and MLP, (2) it enables data reuse/sharing within CHM. The PEs in the same column of CHM can share the same weight block, while the PEs in the same rows of CHM can share the same input token block. This data-sharing strategy simplifies the computation complicated by the irregular data access pattern of block-wise weight pruning. (3) By selecting proper $p_{\text{c}}$ with a load balancing strategy, we can alleviate the load imbalance caused by block-wise weight pruning.
@@ V-C Workflow
The proposed accelerator executes the input model layer-by-layer, with each layer being executed using the same set of computational resources (e.g., MPCA). Each CHM utilizes Sparse Block-wise Matrix Multiplication (SBMM) to process sparse matrices in blocks. For dense matrices, the computation shifts to Dense Block-wise Matrix Multiplication (DBMM) and Dense Head-wise Block Matrix Multiplication (DHBMM) for a focused computation on matrix blocks associated with each head. Each layer performs as follows:
@@ V-C1 MSA Execution
MSA is divided into four stages (shown in Figure 7): stage (i) computes $\mathbf{Q}$, $\mathbf{K}$ and $\mathbf{V}$ through $[\mathbf{Q},\mathbf{K},\mathbf{V}]=\mathbf{Z}[\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}]$. $\mathbf{Z}$ is the input token matrix and [][\;][ ] is matrix concatenation. Let $\mathbf{Q}_{h},\mathbf{K}_{h}$ and $\mathbf{V}_{h}$ denote the query, key, and value matrices for a head $h$, where $(0\leq h<H)$. The algorithm for executing each matrix multiplication (e.g., dense token matrix $\mathbf{Z}$ multiply by sparse weight matrices $\mathbf{W}_{q},\mathbf{W}_{k},\mathbf{W}_{v}$) in MSA using MPCA is shown in Algorithm 2 and an example is shown in Figure 8. Each CHM computes its corresponding head $[\mathbf{Q}_{h},\mathbf{K}_{h},\mathbf{V}_{h}]$. Since we perform block partitioning for each matrix, the computation is executed in a block-wise fashion. Within each CHM, the PEs within a column share the same column of weight, which is stored in the column buffer (CB). PEs of the same row share the same row of $\mathbf{Z}$. We use $\text{PE}(i,j,k)$ to denote a PE in the $k^{\text{th}}$ CHM at location $(i,j)$ within the CHM. A column of PEs: $\text{PE}(:,j,k)$ share the same column of weight with the corresponding header information (See Figure 5). Thus, each PE in $j^{\text{th}}$ column utilizes the shared header indices (Figure 5) to fetch the corresponding data block from the token matrix (in the local input buffer) to perform block-wise matrix multiplication. The partial results are accumulated in local result buffers.
The stage (ii) computes the attention scores $\mathbf{A}_{h}=\mbox{softmax}(\frac{\mathbf{Q}_{h}\mathbf{K_{h}}^{T}}{\sqrt{D^{\prime}}}),\text{ }(0\leq i<H)$. $\mathbf{Q}_{h}\mathbf{K}_{h}^{T}$ is dense block-wise matrix multiplication executed via the MPCA module (See Algorithm 2). $\mathbf{Q}$ is buffered in the GFB, and $\mathbf{K}^{T}$ is buffered in the CB. The output data blocks of $\mathbf{Q}_{h}\mathbf{K}_{h}^{T}$ are sent to EM module for element-wise scaling (by $1/\sqrt{D^{\prime}}$) and exponentiation to obtain $\text{exp}(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{D^{\prime}}})$. Then, we utilize MPCA to compute the scaling factors for $\mbox{softmax}(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{D^{\prime}}})$. The rows of matrix $\text{exp}(\frac{\mathbf{Q}_{h}\mathbf{K}_{h}^{T}}{\sqrt{D^{\prime}}})$ and their corresponding computed scaling factors, are streamed from MPCA to EM to perform the scaling to obtain the attention scores, $\mathbf{A}^{H}$.
@@ V-C2 MLP Execution
1:Input matrix $\mathbf{X}\in\mathbb{R}^{M_{1}\times M_{2}}$; weight matrix $\mathbf{W}=[\mathbf{W}_{0},\mathbf{W}_{1},...,\mathbf{W}_{H-1}]\in\mathbb{R}^{M_{2}\times D}$, where each $\mathbf{W}_{h}\in\mathbb{R}^{M_{2}\times D^{\prime}}$ ($0\leq h<H$); $D=HD^{\prime}$ where $H$ denotes the number of heads and $D^{\prime}$ is the dimension per head; block size $b$
@@
7:// To compute $H$ heads, $p_{h}$ CHMs need $\left\lceil\frac{H}{p_{\text{h}}}\right\rceil$ iterations
8:for $i=0\;\text{ to }\;\left\lceil\frac{H}{p_{\text{h}}}\right\rceil-1$ do
9:    for each $\text{CHM}_{j}$ with $j=0\;\text{ to }\;p_{\text{h}}-1$ Parallel do
10:         // $\text{CHM}_{j}$ computes $\mathbf{Y}_{j+ip_{\text{h}}}$
11:         // To compute $\left\lceil\frac{D^{\prime}}{b}\right\rceil$ column blocks of a $\mathbf{W}_{h}$, $p_{c}$ columns            of PEs in a CHM need $\left\lceil\left\lceil\frac{D^{\prime}}{b}\right\rceil/p_{\text{t}}\right\rceil$ iterations
12:         for $k=0\;\text{ to }\;\left\lceil\left\lceil\frac{D^{\prime}}{b}\right\rceil/{p_{\text{c}}}\right\rceil-1$ do
13:             // Load weights into CB
14:             // To compute $\left\lceil\frac{M}{b}\right\rceil$ row blocks of $\mathbf{X}$, $p_{t}$ rows of PEs                 in a CHM need $\left\lceil\left\lceil\frac{M}{b}\right\rceil/p_{\text{t}}\right\rceil$ iterations.
15:             for $l=0\;\text{ to }\;\left\lceil\left\lceil\frac{M}{b}\right\rceil/p_{\text{t}}\right\rceil-1$ do
16:                 // Load data (partition of $\mathbf{X}$) into GFB
17:                 for each $\text{PE}_{j}(m,n)$ in $\text{CHM}_{j}$ Parallel do
18:                     // $\text{PE}_{j}(m,n)$ computes output block                             $\mathbf{Y}_{j+ip_{\text{h}}}[m+lp_{\text{t}},n+kp_{\text{c}}]$
19:                     if MPCA mode is SBMM then
20:                         Fetch $\mathbf{X}[m+lp_{\text{t}},\text{idx}]$ from GFB for all idx                                in the header of $\mathbf{W}_{j+ip_{\text{h}}}[:,n+kp_{\text{c}}]$
21:                         Compute $\mathbf{Y}_{j+ip_{\text{h}}}[m+lp_{\text{t}},n+kp_{\text{c}}]$ using                                the fetched input blocks from GFB
22:                     else
23:                         // MPCA mode is DBMM
24:                         Fetch $\mathbf{X}[m+lp_{\text{t}},:]$ from GFB
25:                         Compute $\mathbf{Y}_{j+ip_{\text{h}}}[m+lp_{\text{t}},n+kp_{\text{c}}]$ using                                all the input blocks from GFB
@@ V-E1 Resource Consumption Model
We perform theoretical analysis for the performance achieved by the codesign and its hardware resource utilization. We denote the total computational resources utilized by the MPCA, EM and TDHM as $R_{\text{MPCA}},R_{\text{EM}}$ and $R_{\text{TDHM}}$, respectively. $R_{\text{MPCA}}$ is proportional to the total number of computation units: $p_{\text{t}}p_{\text{h}}p_{\text{c}}p_{\text{pe}^{2}}$. Compared to $R_{\text{MPCA}}$, the resources used by $R_{\text{TDHM}}$ and $R_{\text{EM}}$ are negligible, and thus ignored for analysis. The total $R_{\text{Total}}$ (DSPs and LUTs) are $R_{\text{Total}}=(c_{1}p_{\text{t}}p_{\text{h}}p_{\text{c}}p_{\text{pe}}^{2},c_{2}p_{\text{t}}p_{\text{h}}p_{\text{c}}p_{\text{pe}}^{2})$, where $c_{1}$ and $c_{2}$ denote the amount of DSPs and LUTs utilized by a single computation unit. The size of the (global) feature buffer, column buffer and the (global) result buffer, associated with the MPCA, are $b^{2}p_{t}\gamma$, $b^{2}p_{c}\gamma$ and $b^{2}p_{\text{t}}p_{\text{h}}p_{\text{c}}$, respectively. Here, $b$ is the block size and $\gamma$ is a constant that equals the (maximum) total number of row blocks required to compute a single output block. We match the buffer sizes across compute units to improve the dataflow performance of the accelerator. This gives a total buffer size of $4\times\max({b^{2}p_{\text{t}}p_{\text{h}}p_{\text{c}},b^{2}p_{t}\gamma})$ for the EM module and $2\times\max({b^{2}p_{\text{t}}p_{\text{h}}p_{\text{c}},b^{2}p_{t}\gamma})$ for the TDHM module. The EM module requires a buffer to store the input, scaling factor, addtion factor and the output. Similarly, TDHM requires an input and an output buffer. The total size of required buffers $B_{\text{Total}}$ is given as, $B_{\text{Total}}=b^{2}p_{t}\gamma+b^{2}p_{c}\gamma+b^{2}p_{\text{t}}p_{\text{h}}p_{\text{c}}+6\times\max({b^{2}p_{\text{t}}p_{\text{h}}p_{\text{c}},b^{2}p_{t}\gamma})$. $R_{\text{Total}}$ and $B_{\text{Total}}$ are the estimation of resource utilization. The main design $p_{\text{t}}$, $p_{\text{h}}$ and $p_{\text{c}}$, with $\gamma$ are empirically set according the resource of target FPGA platform (See Section VI for details).